A whale A baker purchased 5 lb of wheat flour and 15 lb of rye flour for a total cost of $14.75. A second purchase, at the same prices, included 5 lb of wheat flour and 14 lb of rye flour. The cost of the second purchase was $14.01.
Find the cost per pound of the wheat flour and the rye flour.

Answer by Stitch(470)

Price of Wheat flour per lb = X Price of Rye flour per lb = Y
Equation 1: 5*X + 15*Y = $14.75
Equation 2: 5*X + 14*Y = $14.01
Set both euations equal to zero
Equation 1: 5*X + 15*Y - $14.75 = 0
Equation 2: 5*X + 14*Y - $14.01 = 0
Now we can set the equations equal to each other
5*X + 15*Y - $14.75 = 5*X + 14*Y - $14.01
Subtract 5*X from both sides
15*Y - $14.75 = 14*Y - $14.01
Subtract 14*Y from both sides
Y - $14.75 = -$14.01
Add $14.75 to both sides
Y = $0.74
I plugged $0.74 back in for Y in equation 1. (You can use either equation)
Equation 1: 5*X + 15*Y - $14.75 = 0
5*X + 15*$0.74 - $14.75 = 0
5*X + $11.1 - $14.75 = 0
5*X + -$3.65 = 0
5*X = $3.65
Divide both sides by 5
X = $0.73